# Condensing and Extracting Against Online Adversaries

### v2

@@ 1.3 Limits on Extraction from oNOSF Sources
Our main technique is a generic way to transform leader election and coin flipping protocols (formally defined in Section 3.3) into extractors for oNOBF and oNOSF sources. This is proved in Lemma 8.1; the general idea of constructing an extractor is to simulate an appropriate leader election protocol with the source at hand (oNOBF or oNOSF), and output according to the chosen leader. To instantiate this transformation, we revisit previous leader election protocols in Section 9. Our leader election protocols provide a slightly stronger than usual guarantee: a good player is elected as the leader with probability close to $1$ (see Lemma 9.1 and Lemma 9.5). This contrasts with the usual guarantee in leader election protocols, where a good leader is chosen with only a non-trivial (constant) probability.
1.4 Application to Collective Coin Flipping and Collective Sampling
1.4.1 Background
Say there are $\ell$ players who have a common broadcast channel and want to jointly perform a task such as collectively flipping a coin. Some $b$ players out of them are “bad” and want to deter the task. We assume the bad players are computationally unbounded so cryptographic primitives are of no use. We further assume that each player has private access to uniform randomness. [ben-or_collective_1989] initiated the study of this model and aptly termed this task as “collective coin flipping.”
The simplest way to collectively flip a coin would be for all the players to initially agree on a function $f:\{0,1\}^{\ell}\to\{0,1\}$, then synchronously broadcast one random bit $r_{i}$, and to finally agree on the output being $f(r_{1},\dots,r_{\ell})$. However, synchronizing broadcasts is hard, and it could be that the bad players set their output as function of the bits of the good players. [kahn_influence_1988] showed that no function $f$ can handle more than $O\left(\frac{\ell}{\log\ell}\right)$ corruptions.
One way to allow for more corruptions (almost linear) among players is to consider “protocols” that allow more rounds of communication. In particular, a protocol can be thought of as a tree where each vertex represents a “round” where in every round the following happens: all good players sends their bits, then all bad players send their bits as a function of the bits of the good players, and they jointly compute a function of these bits. Depending on the outcome of the function, everyone branches on one branch in this tree. Furthermore, every leaf is labeled with final outcomes (say 00 or $1$) and, once a leaf is reached, that is the outcome that everybody agrees on. [ggl98] initiated the study of protocols where the outcomes are from a larger range and where the bad players are trying to minimize the largest probability of any outcome. They called this problem “collective sampling.” For a formal definition, see Section 3.3.
1.4.2 Known Results
[ben-or_collective_1989] showed that for protocols with outcomes $\{0,1\}$, $b$ bad players can always ensure that some outcome occurs with probability at least $\frac{1}{2}+\frac{b}{2\ell}$. [alonnoar93collective] first constructed a protocol that can handle a linear number of corruptions. Follow-up works tried to reduce the number of rounds in this protocol where, in some settings, players were allowed to send more than one bit per round [russellzuckerman01, feige_noncryptographic_1999].
[ggl98] showed that for all collective sampling protocols and all outcomes, there exists a way for $b$ bad players to coordinate and ensure that an outcome that happens without corruption with probability $p$, now happens with probability $p^{1-(b/n)}\geq p\left(1+\frac{b}{n}\log(1/p)\right)$. Nearly matching collective sampling protocols were constructed by [ggl98, SV08selection, GVZ06selection]. For an overview of further results and bounds, see [dodis2006fault].
1.4.3 Connection to oNOSF Sources

### v3

@@ 1.3 Extraction from oNOSF Sources
Our main technique for Theorems 5 and 6 is a new generic way to transform leader election and coin flipping protocols (formally defined in Section 4.4) into extractors for oNOBF and oNOSF sources. This is given in Lemma 8.2; the general idea of constructing an extractor is to simulate an appropriate leader election protocol with the source at hand (oNOBF or oNOSF), and output according to the chosen leader. To instantiate this transformation, we revisit previous leader election protocols in Section 9. Our leader election protocols provide a slightly stronger than usual guarantee: a good player is elected as the leader with probability close to $1$ (see Lemma 9.1 and Lemma 9.5). This contrasts with the usual guarantee in leader election protocols, where a good leader is chosen with only a non-trivial (constant) probability. We give more connections to distributed computing in Section 3 where we delineate applications of our results to collection coin flipping and collective sampling.
2 Proof Overview
Our proof overview begins by outlining our new explicit result for condensers in Section 2.1 that is able to handle polylogarithmic block length. Next, we present our transformation of low-entropy oNOSF sources to uniform oNOSF sources in Section 2.2 before discussing our existential results for condensers that can handle constant block length in Section 2.3. We present the main ideas behind our results regarding online influence and extractor lower bounds in Section 2.4. In Section 2.5, we overview our extractor constructions for oNOBF and oNOSF sources that are based on a general transformation from leader election protocols.
2.1 Explicit Condensers for Uniform oNOSF Sources
We sketch here our constructions (as well as proof ideas) of explicit condensers for uniform $\left(0.51\ell,\ell,n\right)$}-oNOSF sources. [^9] Since all sources are uniform here, we will not explicitly mention this again. The goal will be to construct explicit condensers that work with as few good sources as possible while minimizing the block length $n$. In particular, we will show:
Theorem 2.1 (Theorem 5.1, simplified).
For all $0<\varepsilon$ and $n,\ell\in\mathbb{N}$ where $n\geq\left(\frac{\log(\ell)}{\varepsilon}\right)^{\Omega(1)}$, there exists an explicit condenser $\operatorname{\mathsf{2Cond}}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $\left(g=0.51\ell,\ell,n\right)$}-oNOSF source $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-\left(\frac{\log(\ell)}{\varepsilon}\right)^{O(1)}\cdot\log(n)$ where $m=0.001\cdot\ell n-O\left(\ell\log(\ell)\log(1/\varepsilon)\right)$.
We will require two main tools to show this. The first one uses ideas from the leader election literature and allows us to sample a $O(\log(\ell))$ sized committee starting from $\ell$ players while essentially maintaining the fraction of bad players. Formally:
Lemma 2.2 (Lemma 5.3, simplified).
For all $\varepsilon_{s}>0$, $n,\ell\in\mathbb{N}$, and constant $\varepsilon_{a}>0$ where $n\geq\Omega(\log(\ell)\log(1/\varepsilon_{s}))$, there exists an explicit function $\operatorname{\mathsf{oNOSFSamp}}:(\{0,1\}^{n})^{\ell}\to[\ell]^{D}$ where $D\leq O(\log(\ell/\varepsilon_{s}))$ with the following property. [^10] Even though the output domain of $\operatorname{\mathsf{oNOSFSamp}}$ is a vector, we will abuse notation and often treat it as a set. For all $S\subset[\ell]$ and $\left(g,\ell,n\right)$}-oNOSF sources $\mathbf{X}$, we have that
$\Pr_{x\sim\mathbf{X}}\left[\left\lvert\frac{\left\lvert\operatorname{\mathsf{oNOSFSamp}}(x)\cap S\right\rvert}{D}-\frac{\left\lvert S\right\rvert}{\ell}\right\rvert\geq\varepsilon_{a}\right]\leq\varepsilon_{s}$
At a high level, to construct $\operatorname{\mathsf{oNOSFSamp}}$, we slightly modify the committee selection procedure from [RZ01] and instantiate it with a seeded extractor with near optimal dependence on the seed [Zuc07].
Our second tool is a seeded condenser for general min-entropy sources $\mathbf{X}$ that uses an oNOSF source $\mathbf{Y}$ as the seed, where the bad bits in $\mathbf{Y}$ can depend on $\mathbf{X}$.
Lemma 2.3 (Lemma 5.4, simplified).
There exists a constant $C_{\operatorname{\mathsf{2Cond}}}$ such that for all $n_{x},k,n_{y},t\in\mathbb{N}$ with $\varepsilon>0$ and $n_{y}\geq(C_{\operatorname{\mathsf{2Cond}}})^{t}\log(tn_{x}/\varepsilon)$, there exists an explicit condenser $\mathsf{Cond}:\{0,1\}^{n_{x}}\times(\{0,1\}^{n_{y}})^{t}\to\{0,1\}^{m}$ where $m=\frac{1}{3}(k-(C_{\operatorname{\mathsf{2Cond}}})^{t}\log(tn_{x}/\varepsilon))$ so that the following holds: For all $(n_{x},k)$-sources $\mathbf{X}$ and $\left(g=1,\ell=t\right)$}-oNOSF sources $\mathbf{Y}\sim(\{0,1\}^{n_{y}})^{t}$ such that the good blocks in $\mathbf{Y}$ are independent of $\mathbf{X}$ and the bad blocks in $\mathbf{Y}$ can depend on $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\operatorname{\mathsf{2Cond}}(\mathbf{X},\mathbf{Y}))\geq m-(C_{\operatorname{\mathsf{2Cond}}})^{t}\log(tn_{x}/\varepsilon)$.
We will sketch how to construct this in Section 2.1.6. Using these tools, we are ready to present our explicit condenser. In fact, we will provide sketches of five different constructions of explicit condensers with increasingly better parameter dependence; the fifth (and final) one is the construction given by Theorem 2.1. The parameters they will vary in are the fraction of good blocks (i.e., $g/\ell$) and the block length $n$. As we will see, each construction will build on ideas from the previous construction.
2.1.1 Construction 1: 51% good and $n\geq 2^{\Omega(\ell)}$
We here construct the following condenser:
Theorem 2.4 (Theorem 5.6, simplified).
For all $0<\varepsilon$ and $n,\ell\in\mathbb{N}$ where $n\geq 2^{\Omega(\ell)}\log(1/\varepsilon)$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $\left(g=0.51\ell,\ell,n\right)$}-oNOSF source $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-2^{O(\ell)}\log(1/\varepsilon)$ where $m=0.0001\cdot n\ell$.
Proof sketch.
Let $\gamma=0.01$ and $\ell^{\prime}=\ell/2$. We decompose the input $\mathbf{X}$ into two equal sized parts so that $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2})$ where both $\mathbf{X}_{1},\mathbf{X}_{2}\sim(\{0,1\}^{n})^{\ell/2}\equiv(\{0,1\}^{n})^{\ell^{\prime}}$. Since $\mathbf{X}$ has $(0.5+\gamma)\ell$ good players, we infer that each of $\mathbf{X}_{1},\mathbf{X}_{2}$ has at least $\gamma\ell=(2\gamma)\cdot\ell^{\prime}$ good players. In particular, we use the fact that $H_{\infty}(\mathbf{X}_{1})\geq(2\gamma)\cdot(\ell n/2)$ and that $\mathbf{X}_{2}$ is a $\left(g=(2\gamma)\ell/2,\ell/2,n\right)$}-oNOSF source. With this, we let $\operatorname{\mathsf{2Cond}}$ be the condenser from our second tool Lemma 2.3 and let out final output be $\operatorname{\mathsf{2Cond}}(\mathbf{X}_{1},\mathbf{X}_{2})$. Note that here $t=\ell$ and so, this requires $n\geq(C_{\operatorname{\mathsf{2Cond}}})^{\ell}\log(\ell n/\varepsilon)$, an inequality that we indeed satisfy. The guarantees from Lemma 2.3 provide us with the desired claim. ∎
Each of the subsequent constructions will use a similar construction idea as above. However, they will try to decrease $t$ as much as possible, where $t$ is the number of players in the oNOSF source when applying $\operatorname{\mathsf{2Cond}}$ from Lemma 2.3. Note that any reduction results in a decrease in the block length requirement $n$.
2.1.2 Construction 2: 67% good and $n\geq\operatorname{poly}(\ell)$
Our next construction requires a slightly larger fraction of good blocks. However, the block length required is exponentially improved.
Theorem 2.5 (Theorem 5.7, simplified).
For all $0<\varepsilon$ and $n,\ell\in\mathbb{N}$ where $n\geq\left(\frac{\ell}{\varepsilon}\right)^{\Omega(1)}$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $\left(g=0.67\ell,\ell,n\right)$}-oNOSF source $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-\left(\frac{\ell}{\varepsilon}\right)^{O(1)}\log(n)$ where $m=0.001\cdot\ell n$.
Proof sketch.
Let $\gamma=0.67-(2/3)$. Let $\ell^{\prime}=\ell/3$. We decompose the input $\mathbf{X}$ into three equal sized parts so that $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3})$ where all $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\sim(\{0,1\}^{n})^{\ell^{\prime}}$. Since $\mathbf{X}$ has $((2/3)+\gamma)\ell$ good players, we easily see that each $\mathbf{X}_{i}$ is a $\left(g=3\gamma\ell^{\prime},\ell^{\prime},n\right)$}-oNOSF source.
We use $\mathbf{X}_{1}$ to sample a $O(\log(\ell))$ sized committee from $\mathbf{X}_{3}$. To do so, we use $\operatorname{\mathsf{oNOSFSamp}}_{1\to 3}:(\{0,1\}^{n})^{\ell^{\prime}}\to(\{0,1\}^{\ell^{\prime}})^{D}$ from Lemma 2.2 with $S\subset[\ell^{\prime}]$ being the set of good players from $\mathbf{X}_{3}$, the approximation factor $\varepsilon_{a}=\gamma$ and sampling error $\varepsilon_{s}=\varepsilon/3$. Let $\mathcal{C}_{3}\subset[\ell^{\prime}],\left\lvert\mathcal{C}_{3}\right\rvert\leq D_{3}=O(\log(\ell^{\prime}/\varepsilon))=O(\log(\ell/\varepsilon))$ be the committee of players thus obtained. The approximation property of the sampler guarantees us that out of $\geq 3\gamma\ell^{\prime}$ good players in $\mathbf{X}_{3}$, at least $2\gamma\left\lvert\mathcal{C}_{3}\right\rvert$ many good players will be in $\mathcal{C}_{3}$ with probability $1-\varepsilon/3$. Let $\mathbf{Y}_{3}$ be the $\left(2\gamma D_{3},D_{3},n\right)$}-oNOSF source obtained by restricting the players in $\mathbf{X}_{3}$ to the committee $\mathcal{C}_{3}$. We finally use $\operatorname{\mathsf{2Cond}}$ from Lemma 2.3 and output $\operatorname{\mathsf{2Cond}}(\mathbf{X}_{2},\mathbf{Y}_{3})$. Here, the parameter $t$ in Lemma 2.3 will be set to $D\leq O(\log(\ell/\varepsilon))$, and hence, Lemma 2.3 would only require that $n\geq(C_{2_{\operatorname{\mathsf{2Cond}}}})^{t}\log(\ell n/\varepsilon)=\left(\frac{\ell}{\varepsilon}\right)^{O(1)}$, a condition that we do meet. We carefully compute the remaining parameters to infer the claim. ∎
2.1.3 Construction 3: 76% good and $n\geq\operatorname{poly}(\log(\ell))$
We build on our previous construction and show how to condense when the block length requirement is again exponentially decreased. This comes at a cost of slightly larger fraction of good blocks.
Theorem 2.6 (Theorem 5.9, simplified).
For all $0<\varepsilon$ and $n,\ell\in\mathbb{N}$ where $n\geq\left(\frac{\log(\ell)}{\varepsilon}\right)^{\Omega(1)}$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $\left(g=0.76\ell,\ell,n\right)$}-oNOSF source $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-\left(\frac{\log(\ell)}{\varepsilon}\right)^{O(1)}\log(n)$ where $m=0.001\cdot\ell n$.
Proof sketch.
Let $\gamma=0.01$. Let $\ell^{\prime}=\ell/4$. We decompose the input $\mathbf{X}$ into four equal sized parts such that $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3},\mathbf{X}_{4})$ and conclude that each $\mathbf{X}_{i}$ is a $\left(g=4\gamma\ell^{\prime},\ell^{\prime},n\right)$}-oNOSF source. However here, we claim something stronger. Call $i\in[\ell^{\prime}]$ a totally good index if it corresponds to a good player across each of the four blocks. Since $\mathbf{X}$ has $(3/4+\gamma)\cdot(4\ell^{\prime})$ good players, there must be $\geq 4\gamma\ell^{\prime}$ totally good indices.
We first use $\mathbf{X}_{1}$ to sample a $D_{2}=O(\log(\ell/\varepsilon))$ sized committee $\mathcal{C}_{2}\subset[\ell^{\prime}]$ using the sampler from Lemma 2.2 such that $\mathcal{C}_{2}$ will have at least $3\gamma$ fraction of totally good indices. We let $\mathbf{Y}_{2}$ be the $\left(g=3\gamma D_{2},D_{2},n\right)$}-oNOSF source obtained by restricting $\mathbf{X}_{2}$ to indices from $\mathcal{C}_{2}$.
Second, we use $\mathbf{Y}_{2}$ to sample a $D_{4}=O(\log(\log(\ell)/\varepsilon))$ sized committee $\mathcal{C}_{4}\subset\mathcal{C}_{2}$ such that $\mathcal{C}_{4}$ has at least $2\gamma$ fraction of totally good indices. We let $\mathbf{Y}_{4}$ be the $\left(g=2\gamma D_{4},D_{4},n\right)$}-oNOSF source obtained by restricting $\mathbf{X}_{4}$ to indices from $\mathcal{C}_{4}$.
Third and last, we use $\operatorname{\mathsf{2Cond}}$ from Lemma 2.3 and output $\operatorname{\mathsf{2Cond}}(\mathbf{X}_{3},\mathbf{Y}_{4})$. Here, the parameter $t$ in Lemma 2.3 will be set to $D_{4}\leq O(\log(\log(\ell)/\varepsilon))$, and hence, Lemma 2.3 would only require that $n\geq(C_{2_{\operatorname{\mathsf{2Cond}}}})^{t}\log(\ell n/\varepsilon)=\left(\frac{\log(\ell)}{\varepsilon}\right)^{O(1)}$, a condition that we do meet. We carefully compute the remaining parameters to infer the claim. ∎
2.1.4 Construction 4: 67% good and $n\geq\operatorname{poly}(\log(\ell))$
Our next construction maintains a similar guarantee as before on the block length and decreases the requirement on the fraction of good blocks.
Theorem 2.7 (Theorem 5.10, simplified).
For all $0<\varepsilon$ and $n,\ell\in\mathbb{N}$ where $n\geq\left(\frac{\log(\ell)}{\varepsilon}\right)^{\Omega(1)}$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any $\left(g=0.67\ell,\ell,n\right)$}-oNOSF source $\mathbf{X}$, we have that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-\left(\frac{\log(\ell)}{\varepsilon}\right)^{O(1)}\log(n)$ where $m=0.001\cdot\ell n$.
Proof sketch.
Let $\gamma=0.67-(2/3)$. Let $\ell^{\prime}=\ell/3$. We decompose $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3})$ so that each $\mathbf{X}_{i}$ is a $\left(g=3\gamma\ell^{\prime},\ell^{\prime},n\right)$}-oNOSF source with $3\gamma\ell^{\prime}$ totally good indices.
For the first step, we again use $\mathbf{X}_{1}$ along with a sampler from Lemma 2.2 to obtain $\mathcal{C}_{2}\subset[\ell^{\prime}]$ with $D_{2}=\left\lvert\mathcal{C}_{2}\right\rvert\leq O(\log(\ell/\varepsilon)$ and a subsource of $\mathbf{X}_{2}$ restricted to $\mathcal{C}_{2}$ - $Y_{2}$ that is a $\left(g=2\gamma D_{2},D_{2},n\right)$}-oNOSF source. For the second step, we again use $\mathbf{Y}_{2}$ to sample a $D_{3}=O(\log(\log(\ell)/\varepsilon))$ sized committee $\mathcal{C}_{3}\subset\mathcal{C}_{2}$ such that $\mathcal{C}_{3}$ has $\geq\gamma$ fraction of totally good indices. We let $\mathbf{Y}_{3}$ be the $\left(g=\gamma D_{3},D_{3},n\right)$}-oNOSF source obtained by restricting $\mathbf{X}_{3}$ to indices from $\mathcal{C}_{3}$. Third and last, we use $\operatorname{\mathsf{2Cond}}$ from Lemma 2.3 and output $\operatorname{\mathsf{2Cond}}(\mathbf{X}_{2},\mathbf{Y}_{3})$. Again the parameter $t$ in Lemma 2.3 will be set to $D_{3}\leq O(\log(\log(\ell)/\varepsilon))$ and would only require that $n\geq\left(\frac{\log(\ell)}{\varepsilon}\right)^{O(1)}$.
Analyzing this construction requires more care since we use $\mathbf{X}_{2}$ to both sample from $\mathbf{X}_{3}$ and as a source for $\operatorname{\mathsf{2Cond}}$. We use the chain rule for min-entropy (Lemma 4.3) to argue that most fixings of $\mathbf{Y}_{2}=y_{2}$ will leave $\mathbf{X}_{2}$ with lot of entropy (since sampler requires few random bits) and observe that such a fixing still leaves $\mathbf{X}_{3}$ as oNOSF source with the same parameters. Also since for most fixings of $\mathbf{Y}_{2}=y_{2}$, the committee $\mathcal{C}_{3}$ has $\gamma$ fraction of good players, we obtain our claim. ∎
2.1.5 Construction 5: 51% good and $n\geq\operatorname{poly}(\log(\ell))$
We lastly construct the condenser promised in our main result - Theorem 2.1.
Proof sketch of Theorem 2.1.
Let $\gamma=0.01$. Let $\ell^{\prime}=\ell/2$. We decompose $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2})$ so that each $\mathbf{X}_{i}$ is a $\left(g=2\gamma\ell^{\prime},\ell^{\prime},n\right)$}-oNOSF source with $2\gamma\ell^{\prime}$ totally good indices.
For the first step, we let $\mathbf{X}_{1}^{\prime}$ be the $\left(2\gamma\ell^{\prime},\ell^{\prime},n_{1}^{\prime}\right)$}-oNOSF source obtained by taking a prefix of length $n_{1}^{\prime}$ from each source where we set $n_{1}^{\prime}\ll n$ but also $n_{1}^{\prime}$ is long enough to be used by the sampler from Lemma 2.2. We use $\mathbf{X}_{1}^{\prime}$ with such a sampler to obtain a committee $\mathcal{C}_{1\to 2}\subset[\ell^{\prime}]$ with $D_{1\to 2}=\left\lvert\mathcal{C}_{1\to 2}\right\rvert\leq O(\log(\ell/\varepsilon)$ and a subsource of $\mathbf{X}_{2}$ restricted to $\mathcal{C}_{1\to 2}$, which we call $\mathbf{Y}_{1\to 2}$, that is a $\left(g=(3/2)\gamma D_{1\to 2},D_{1\to 2},n\right)$}-oNOSF source.
We argue that: 1) most fixings of $\mathbf{X}_{1}^{\prime}$ are such that they leave $\mathbf{X}_{1}$ with high entropy, and 2) that the committee $\mathcal{C}_{1\to 2}$ obtained will have $3\gamma/2$ fraction of good players. We obtain this using the chain rule for min-entropy and by guarantees of the sampler. We condition on such a fixing from here on.
For the second step, we use $\mathbf{Y}_{1\to 2}$ to sample a $D_{2\to 2}=O(\log(\log(\ell)/\varepsilon))$ sized committee $\mathcal{C}_{2\to 2}\subset\mathcal{C}_{1\to 2}$ such that $\mathcal{C}_{2\to 2}$ has $\geq\gamma$ fraction of totally good indices. We let $\mathbf{Y}_{2\to 2}$ be the $\left(g=\gamma D_{2\to 2},D_{2\to 2},n\right)$}-oNOSF source obtained by restricting $\mathbf{X}_{2}$ to indices from $\mathcal{C}_{2\to 2}$.
Lastly, we use $\operatorname{\mathsf{2Cond}}$ from Lemma 2.3 and output $\operatorname{\mathsf{2Cond}}(\mathbf{X}_{1},\mathbf{Y}_{2\to 2})$. The parameter $t$ in Lemma 2.3 will be set so that it would only require that $n\geq\left(\frac{\log(\ell)}{\varepsilon}\right)^{O(1)}$. We must be a bit careful about the error parameter $\varepsilon_{\operatorname{\mathsf{2Cond}}}$ for $\operatorname{\mathsf{2Cond}}$ and we will choose it to be extremely small.
Analyzing this construction requires care since we use $\mathbf{X}_{2}$ both to sample from within $\mathbf{X}_{2}$ itself and also as a seed for $\operatorname{\mathsf{2Cond}}$. We cannot use the chain rule since a fixing of $\mathbf{Y}_{1\to 2}$ will destroy the structure of the source $\mathbf{X}_{2}$. We first see that the sampler guarantees that no matter how the adversary behaves, with probability $1-\varepsilon/4$, the sampler will succeed in selecting a committee $\mathcal{C}_{2\to 2}$ with $\gamma$ fraction of good players. We pay $\varepsilon/4$ in error and now assume that the sampler always succeeds in doing so. We then compare two scenarios: one scenario $\mathsf{Opt}$ where the bits in $\mathbf{Y}_{1\to 2}$ are all uniform and independent of all other bits and another scenario $\mathsf{Adv}$ where the bits in $\mathbf{Y}_{1\to 2}$ are all controlled by an adversary (guarantees on this latter scenario suffice for our claim). Let the total number of bits in $\mathbf{Y}_{1\to 2}$ be equal to $b_{1\to 2}$. Under scenario $\mathsf{Opt}$, we easily see that we succeed and with error $\varepsilon_{\operatorname{\mathsf{2Cond}}}$ will have high entropy - say $k$. To compare this to scenario $\mathsf{Adv}$, we use Lemma 7.9 that lets us conclude that in scenario $\mathsf{Adv}$, with error $\varepsilon_{\operatorname{\mathsf{2Cond}}}\cdot 2^{b_{1\to 2}}$, the output will have entropy $k-b_{1\to 2}$. Since we carefully chose $\varepsilon_{\operatorname{\mathsf{2Cond}}}$ to be small enough and $b_{1\to 2}$ is small since we only use $\mathbf{Y}_{1\to 2}$ as source for the sampler, the output will still have small error and will have high-entropy as desired. ∎
2.1.6 Construction of $\operatorname{\mathsf{2Cond}}$
We now sketch how to construct our desired $\operatorname{\mathsf{2Cond}}$.
Proof sketch of Lemma 2.3.
Let $\mathbf{X}$ and $\mathbf{Y}=(\mathbf{Y}_{1},\dots,\mathbf{Y}_{t})$ be the two sources. For $i\in[t]$, let $n_{i}\approx C^{t-i+1}\log(tn_{x}/\varepsilon)$ where $C$ is a large constant. For $i\in[t]$, let $\mathbf{Z}_{i}$ be the length $n_{i}$ prefix of the block $\mathbf{Y}_{i}$. Our final construction will be the parity of the outputs of seeded extractors applied with source $\mathbf{X}$ and seeds $\mathbf{Z}_{i}$. More formally, we output
$\bigoplus_{i=1}^{t}\mathsf{sExt}_{i}(\mathbf{X},\mathbf{Z}_{i}),$
We proceed to sketch the analysis. We are guaranteed that there exists at least one $j\in[t]$ from $\mathbf{Y}$ that is good. We first condition on fixing blocks $\mathbf{Z}_{1},\dots,\mathbf{Z}_{j-1}$. Since these blocks can depend on $\mathbf{X}$, we apply the chain rule for min-entropy (Lemma 4.3) and conclude $\mathbf{X}$ will only lose some small amounts of entropy (the amount will be very small since these blocks are tiny compared to the amount entropy in $\mathbf{X}$). Moreover, since the adversary is online, $\mathbf{Z}_{j}$ remains uniform even after this fixing. We now view our construction as
$g(\mathbf{X})\oplus\bigoplus_{i=j}^{t}\mathsf{sExt}_{i}(\mathbf{X},\mathbf{Z}_{i})$
We now compare two scenarios: (1) Where all of $\mathbf{Z}_{j},\dots,\mathbf{Z}_{t}$ are uniform (2) Only $\mathbf{Z}_{j}$ is uniform and $\mathbf{Z}_{j+1},\dots,\mathbf{Z}_{t}$ are arbitrarily controlled by an adversary and can even depend on $\mathbf{X}$:
In the first scenario, we further condition on fixing $\mathbf{Z}_{j+1},\dots,\mathbf{Z}_{t}$. Since in this scenario $\mathbf{Z}_{i}$ are independent and random, $\mathbf{X}$ retains the same entropy and $\mathbf{Y}_{j}$ remains uniform. So our overall output is of the form $h(\mathbf{X})\oplus\mathsf{sExt}_{j}(\mathbf{X},\mathbf{Z}_{j})$ for some fixed function $h$. We condition on fixing output $h(\mathbf{X})$. Since the number of output bits $m\ll H_{\infty}(\mathbf{X})$, we apply the chain rule to infer that $\mathbf{X}$ still has lots of entropy when we do this fixing. Now the output is just $z\oplus\mathsf{sExt}_{j}(\mathbf{X},\mathbf{Z}_{j})$ where $z$ is a fixed string, and hence is uniform.
The second scenario is more realistic and, in the worst case, this is what can actually happen. We then use the result that if an adversary controls few bits in the input distribution, then they cannot make the output of the condenser too bad (see Lemma 7.9 for the full statement). With this, since we carefully chose geometrically decreasing lengths of $\mathbf{Z}_{i}$ to help control the error, we indeed obtain that the output will be condensed. ∎
2.2 Converting Low-Entropy oNOSF Sources to Uniform oNOSF Sources
They key part of our proof for condensing from low-entropy oNOSF sources is a transformation from low-entropy oNOSF sources to uniform oNOSF sources . Here, we sketch the proof for our transformation in Theorem 1.4 and compare it to that of [CGR24]. Both these transformations rely on two-source extractors (see Definition 4.7 for definition) as a basic primitive.
Given a $\left(g,\ell,n,k\right)$}-oNOSF source $\mathbf{X}=\mathbf{X}_{1},\dots,\mathbf{X}_{\ell}$, [CGR24] uses excellent existential two-source extractors (such as from Lemma 6.6) to define output blocks $\mathbf{O}_{i}=2\mathsf{Ext}(\mathbf{X}_{1}\circ\cdots\circ\mathbf{X}_{i-1},\mathbf{X}_{i})$ for $i\in\{2,\dots,\ell\}$ and define their transformation as $f(\mathbf{X})=\mathbf{O}_{2},\dots,\mathbf{O}_{\ell}$. They show that $\mathbf{O}_{i}$ is a good block if: (1) $\mathbf{X}_{i}$ is a good block and (2) at least one block amongst $\mathbf{X}_{1},\dots,\mathbf{X}_{i-1}$ is a good block. They showed that such a good block will be uniform and independent of the blocks $\mathbf{O}_{2},\dots,\mathbf{O}_{i-1}$ and argued there will be g1g-1 such good output blocks. This indeed shows their output is a uniform $\left(g-1,\ell-1,m\right)$}-oNOSF source. However, each of their output blocks has length $m=O\left(\frac{k}{\ell}\right)\leq O\left(\frac{n}{\ell}\right)$, and so they were not able to handle the case of $n=o(\ell)$. We improve on their construction by using a “sliding window” based technique to obtain a much better transformation that can even handle $n=\operatorname{poly}(\log(\ell))$.
Theorem 2.8 (Theorem 6.1 restated).
The parameter $d$ in our theorem statement above is the width of our sliding window. When we set $d=\ell$ we recover the analysis of [CGR24]. The true advantage of our transformation emerges when $d$ is very small compared to $\ell$. For instance, when $g=0.51\ell,n=\operatorname{poly}(\log(\ell))$ and $k=\operatorname{poly}(\log(\ell))$, we set $d$ to be a large constant and conclude that the output distribution is a uniform $\left(0.509\ell,\ell,\operatorname{poly}(\log(\ell)\right)$}-NOSF source.
Proof sketch of Theorem 2.8.
We first compute the number of good output blocks $g_{out}$. Let $j_{1},\dots,j_{g}$ be the indices of the good input blocks in $\mathbf{X}$ and $d_{i}=j_{i+1}-j_{i}$ be the gap between the $i$-th good block and the next ($i+1$)-th good block. If the gap $d_{i}$ is at most $d$, then $\mathbf{O}_{i+1}$ must be a good output block. So, $g_{out}$ is the number of $i$ such that $d_{i}\leq d$. Since $g\geq 0.51\ell$, such large gaps can’t appear too often and we compute that $g_{out}\geq g-\frac{\ell-g+2}{d}$ as desired.